# Improving Antibody Humanness Prediction using Patent Data

### v1

@@ 4.2 Empirical Analysis of the Pre-training Stage
Figure 2 shows the clustering of sequences from test set of the PAD, based on their origin and potential targets (Figure 2a-b). We see that the sequences from same species as well as those associated with same potential targets tend to cluster together. Note that when training our models, sequence embeddings are generated by aggregating over the amino acids to prevent the model from memorizing the properties that depend on the sequence length (though it might be useful to learn this dependency in some context).
@@ 4.3 Empirical Analysis of the Fine-tuning Stage
A humanness score should be able to distinguish between human antibodies and antibodies from other species. In particular, it should enable doing so for therapeutic antibodies since those are the primary subjects of humanness analysis. The goal of humanness evaluation is to capture and reduce the immunogenicity risk. If antibodies are not structurally similar to natural human antibodies, there is an increased risk of the immune system recognizing them as foreign and producing antibodies against therapeutic antibodies i.e. ”anti-drug antibodies” (ADAs) or ”immunogenicity”. Therefore, we fine-tune and evaluate SelfPAD to predict the humanness of sequences. We refer to the fine-tuned model as FT-SelfPAD. We should note that, in the machine learning literature, the term ”fine-tuning” is used to mean supervised training of a model, which is pre-trained on a large corpus of data (typically in unsupervised framework), on a small dataset. In our context, fine-tuning corresponds to further training of pre-trained model on the same dataset, but using a supervised loss instead of a contrastive loss so that the model can learn to distinguish between human and nonhuman sequences. Once trained, we benchmark FT-SelfPAD against three strong baselines, namely Hu-mAb (Marks et al., 2021), OASis (Prihoda et al., 2022) and AbNatiV (Ramon et al., 2024), by using four datasets: i) 29k unpaired sequences from the test set of the PAD, ii / iii) paired sequences from 553 therapeutics and 217 immunogenicity data (Prihoda et al., 2022) and iv) paired sequences from 25 humanization data (Marks et al., 2021).
@@
29k sequences from the test set of PAD Once the model is fine-tuned for humanness prediction, we extract embeddings of the test set, and cluster sequences based on their chain type and origin. As shown in Figure 2c, the sequences from different origins and chain types make up distinct clusters. For examples, heavy chains from humans (orange in Figure 2c) cluster together, but are separate from the heavy chains of other species (purple). Similarly, nanobodies (i.e. VHH sequences) constitute a distinct cluster away from the rest of the sequences. This result indicates that our model should be able to distinguish the origin and chain type of a sequence. Moreover, we note that lambda chains from both humans and other species cluster together, hinting that our model might have more difficulty in distinguishing the origin of lambda chains. We also plot sequences based on their origin by using PCA of embeddings in Figure 3a, in which we clearly see that FT-SelfPAD can distinguish the sequences from humans, alpaca and other species. Table 2 summarizes the performance of FT-SelfPAD and other baselines for humanness prediction and shows that FT-SelfPAD performs the best across all metrics. We should note that AbNatiV (Ramon et al., 2024) has four models, one for each chain type of heavy, lambda, kappa and VHH. Thus, we used AbNatiV’s corresponding models to predict the humanness (with a threshold of 0.8) of heavy, kappa and lambda chains from human and other species in the test set after removing VHH sequences. For OASis, we evaluate and compare humanness by using the OASis identity score and prevalence threshold of ’relaxed’. We evaluated OASis at three different OASis identity threshold: 0.5, 0.7 and 0.8. We obtained the best result by using identity threshold of 0.7, which we also use in the rest of the experiments. Finally, Hu-mAb relies on separate classifiers, each of which is trained for each human V gene type. It also computes humanness score based on each chain type i.e. heavy or light chain. If it identifies any V gene in the sequence as human, we consider the entire sequence as human. Hence, we scored heavy and light chains from humans and other species after removing VHH sequences. In our experiments, we noticed that Hu-mAb performs poorly on light chains from humans, an observation that is also reported in (Marks et al., 2021), resulting in a poor performance overall.
@@
553 therapeutics data We process the dataset such that 197 human and 226 humanized sequences are labeled as human while the rest is labelled as nonhuman with the assumption that humanized sequences is more likely to be similar to human sequences. We evaluate the models under two scenarios: i) Paired: We measure the humanness prediction of a paired sequence. If at least one of two chains (light and heavy) in a paired sequence is predicted as human, we consider the paired sequence as human. ii) Single chain: We measure the performance of models for each chain. Table 3 summarizes the performance of all models, showing that FT-SelfPAD performs the best for both scenarios.
217 immunogenicity data It contains reported ADA responses from 217 sequences. Similar to (Prihoda et al., 2022), we label sequences with immunogenicity $\geq 10$ as nonhuman while the ones with $<10$ as human. However, we should emphasize that there is a difference between humanness and immunogenicity, and our evaluation of humanness is only a proxy in this context. For example, some fully human antibodies given at high dosage for extended periods of time can end up being more immunogenic than humanised antibodies administered once at low dose. Keeping this in mind, we list our results in Table 4, showing that FT-SelfPAD gives the best performance for the most important metrics, namely ROC AUC and Precision-Recall (PR) AUC, for paired sequences while it performs best across all metrics, except PR-AUC, for the single chains.
25 humanization data The data consists of 25 parental sequences with immunogenicity issues and their corresponding humanized versions. We label parental sequences as nonhuman while humanized ones as human. As shown in Table 5, FT-SelfPAD classifies all sequences correctly. Moreover, we plot both parental sequences and humanized counterparts by using a PCA on the their embeddings from FT-SelfPAD as shown in Figure 3d. We also overlay the sequences from the test set of PAD to give some context. By comparing to the Figure 3a, we see that parental sequences show up in clusters of other species while the humanized sequences appear among the sequences from human origin. Hence, our model is correctly able to identify the sequences with immunogenicity issues. We made similar comparisons for both 553 therapeutics and 217 immunogenicity data in Figure 3b-c, resulting in a similar observation.
@@
Interpretability To enhance the model’s utility in antibody engineering, humanness evaluation should offer interpretable scores by highlighting the impact of individual residues or sequence segments. Improved interpretability includes providing explanations of humanness scores through comparisons with closely related reference sequences. We tested FT-SelfPAD’s ability to infer positions for humanizing sequences by analyzing attention in 25 humanization data sequences. To identify potential candidates for humanization, we subtracted attention scores (extracted from the last layer of encoder) of humanized sequences from their corresponding parental sequences, scaled the results to $[0,1]$, and applied a threshold of $0.7$ for binarization. Figure 6 displays positions with the highest attention, comparing inferred positions with mutated positions in the humanized sequence of Certolizumab (see Section C in the appendix for more results). The model focuses on positions initially mutated in the parental sequence, recommending fewer mutations than those found in the humanized sequences. This aligns with observations in (Marks et al., 2021), suggesting the potential for fewer mutations in humanization.
@@
Ablation study We performed an ablation study on the 217 immunogenicity dataset to analyze design choices in FT-SelfPAD. Table 6 shows that incorporating shift and swap noise, along with re-initializing four pre-trained model layers before fine-tuning, improves performance. Notably, the model is less effective when trained from scratch or fine-tuned without random initialization of its four layers (see first and second rows in Table 6).
@@ 5 Conclusion
In this work, we present a framework, SelfPAD, to learn representations from the Patented Antibody Database, and demonstrate its usefulness by providing a robust model for humanness prediction - a practical challenge in therapeutic antibody development. SelfPAD opens new avenues for advancing the understanding and optimization of antibody sequences, ultimately facilitating the development of more effective and safe therapeutic interventions. Despite its advantages, one of the main drawbacks of our approach is that the PAD is a noisy dataset and further curation is needed to improve the signal-to-noise ratio. Moreover, the PAD introduces new biases e.g. it biases towards lab optimised rather than natural sequences. Thus, we can also improve our results by incorporating data from other sources such as the OAS since it includes a large number of natural sequences. Finally, although this work focuses on humanness prediction, we can fine-tune SelfPAD for other downstream tasks such as predicting developability properties of antibodies.

### v3

@@ 4.2 Empirical Analysis of the Pre-training Stage
Figure 2 shows a clustering of sequences from the test set of the PAD, based on their origin and potential targets (Panels 2a-b). We see that the sequences from same species as well as those associated with same potential targets tend to cluster together. Note that when training our models, sequence embeddings are generated by aggregating over the amino acids to prevent the model from memorizing the properties that depend on the sequence length (though it might be useful to learn this dependency in some context).
@@ 4.3 Empirical Analysis of the Fine-tuning Stage
A humanness score should be able to distinguish between human antibodies and antibodies from other species. In particular, it should enable doing so for therapeutic antibodies since those are the primary subjects of humanness analysis. The goal of humanness evaluation is to capture and reduce the immunogenicity risk. If antibodies are not structurally similar to natural human antibodies, there is an increased risk of the immune system recognizing them as foreign and producing antibodies against therapeutic antibodies, i.e., ”anti-drug antibodies” (ADAs) or ”immunogenicity”. Therefore, we fine-tune and evaluate SelfPAD to predict the humanness of sequences. We refer to the fine-tuned model as FT-SelfPAD. We should note that, in the machine learning literature, the term fine-tuning typically refers to supervised training of a model, which is pre-trained on a large corpus of data (typically using an unsupervised learning algorithm), on a small dataset. In our context, fine-tuning corresponds to further training of pre-trained model on the same dataset, but using a supervised loss instead of a contrastive loss so that the model can learn to distinguish between human and nonhuman sequences. Once trained, we benchmark FT-SelfPAD against three strong baselines, namely Hu-mAb (Marks et al., 2021), OASis (Prihoda et al., 2022) and AbNatiV (Ramon et al., 2024), by using four datasets: i) 29k unpaired sequences from the test set of the PAD, ii / iii) paired sequences from 553 therapeutics and 217 immunogenicity data (Prihoda et al., 2022) and iv) paired sequences from 25 humanization data (Marks et al., 2021).
@@
29k sequences from the test set of PAD. Once the model is fine-tuned for humanness prediction, we extract embeddings of the test set, and cluster sequences based on their chain type and origin. As shown in Figure 2c, the sequences from different origins and chain types make up distinct clusters. For examples, heavy chains from humans (orange in Figure 2c) cluster together, but are separate from the heavy chains of other species (purple). Similarly, nanobodies (i.e., VHH sequences) constitute a distinct cluster away from the other sequences. This result indicates that our model should be able to distinguish the origin and chain type of a sequence. Moreover, we note that lambda chains from both humans and other species cluster together, hinting that our model might have more difficulty in distinguishing the origin of lambda chains. We also plot sequences based on their origin by using PCA of embeddings in Figure 3a, in which we clearly see that FT-SelfPAD can distinguish the sequences from humans, alpaca and other species. Table 2 summarizes the performance of FT-SelfPAD and other baselines for humanness prediction and shows that FT-SelfPAD performs the best across all the metrics. We should note that AbNatiV (Ramon et al., 2024) has four models, one for each chain type of heavy, lambda, kappa and VHH. Thus, we used AbNatiV’s corresponding models to predict the humanness (with a threshold of 0.8) of heavy, kappa and lambda chains from human and other species in the test set after removing VHH sequences. For OASis, we evaluate and compare humanness by using the OASis identity score and prevalence threshold of ’relaxed’. We evaluated OASis at three different OASis identity threshold: 0.5, 0.7 and 0.8. We obtained the best result by using identity threshold of 0.7, which we also use in the remaining experiments. Finally, Hu-mAb relies on separate classifiers, each of which is trained for each human V gene type. It also computes humanness score based on each chain type, i.e., heavy or light chain. If it identifies any V gene in the sequence as human, we consider the entire sequence as human. Hence, we scored heavy and light chains from humans and other species after removing VHH sequences. In our experiments, we noticed that Hu-mAb performs poorly on light chains from humans, an observation that is also reported in (Marks et al., 2021), resulting in a poor performance overall.
@@
553 therapeutics data. We process the dataset such that 197 human and 226 humanized sequences are labeled as human while the others are labelled as nonhuman with the assumption that humanized sequences are more likely to be similar to human sequences. We evaluate the models under two scenarios: i) Paired: We measure the humanness prediction of a paired sequence. If at least one of two chains (light and heavy) in a paired sequence is predicted as human, we consider the paired sequence as human. ii) Single chain: We measure the performance of models for each chain. Table 3 summarizes the performance of all models, showing that FT-SelfPAD performs the best for both scenarios.
217 immunogenicity data. The dataset contains reported ADA responses from 217 sequences. Similar to Prihoda et al. (2022), we label sequences with immunogenicity $\geq 10$ as nonhuman while the ones with $<10$ as human. However, we should emphasize that there is a difference between humanness and immunogenicity, and our evaluation of humanness is only a proxy in this context. For example, some fully human antibodies given at high dosage for extended periods of time can end up being more immunogenic than humanised antibodies administered once at low dose. Keeping this in mind, we summarize our results in Table 4, showing that FT-SelfPAD gives the best performance for the most important metrics, namely ROC AUC and Precision-Recall (PR) AUC for paired sequences—it generally performs the best across all metrics, except PR-AUC for the single chains.
25 humanization data. The dataset consists of 25 parental sequences with immunogenicity issues and their corresponding humanized versions. We label parental sequences as nonhuman and humanized ones as human. As shown in Table 5, FT-SelfPAD classifies all sequences correctly. Moreover, we plot both parental sequences and humanized counterparts by using a PCA on the their embeddings from FT-SelfPAD as shown in Figure 3d. We also overlay the sequences from the test set of PAD to give some context. By comparing to the Figure 3a, we see that parental sequences show up in clusters of other species while the humanized sequences appear among the sequences from human origin. Hence, it appears that our model is able to correctly identify the sequences with immunogenicity issues. We made similar comparisons for both 553 therapeutics and 217 immunogenicity data in Figure 3b-c, resulting in a similar observation.
@@
Interpretability. To enhance the model’s utility in antibody engineering, humanness evaluation should offer interpretable scores by highlighting the impact of individual residues or sequence segments. Improved interpretability includes providing explanations of humanness scores through comparisons with closely related reference sequences. We tested FT-SelfPAD’s ability to infer positions for humanizing sequences by analyzing attention in 25 humanization data sequences. To identify potential candidates for humanization, we subtracted attention scores (extracted from the last layer of encoder) of humanized sequences from their corresponding parental sequences, scaled the results to $[0,1]$, and applied a threshold of $0.7$ for binarization. Figure 4 displays positions with the highest attention, comparing inferred positions with mutated positions in the humanized sequence of Certolizumab (see Section C in the appendix for more results). The model focuses on positions initially mutated in the parental sequence, recommending fewer mutations than those found in the humanized sequences. This aligns with observations in (Marks et al., 2021), suggesting the potential for fewer mutations in humanization.
Alignment with biological principles. To understand the model’s alignment with biological principles, we compared sequences from the PAD’s test set and the 217 Immunogenicity dataset, and analyzed global relationships and differences among closely related sequences. Please use Figure 3a and  3c as reference for the following analysis.
First, we extracted PCA embeddings for human sequences in the PAD test set (orange cluster in Figure 3a), identifying 200 central sequences—90 light chains and 110 heavy chains. These central sequences were used as reference points. Next, we extracted PCA embeddings for sequences in the 217 Immunogenicity dataset (Figure 3c). We then aligned the heavy chains and separately aligned the light chains from both datasets using the Kabat numbering scheme. This allowed us to compare sequence similarities across datasets. We calculated maximum and average sequence similarities between correctly classified, misclassified, false positive, and false negative sequences in the 217 dataset and the central sequences from the PAD test set. Table 6 summarizes this analysis with the maximum and average similarities between sequences in the 217 dataset and the central sequences from the PAD test set. As we move from left to right in the table, the sequence similarity decreases, illustrating global sequence relationships.
Our findings show that correctly predicted sequences have higher maximum and average similarity to the central sequences compared to the misclassified ones. False positives also exhibit higher similarities than false negatives. This demonstrates a meaningful correlation between a sequence’s Euclidean distance from the center of the human cluster in PCA space and its sequence similarity to the central sequences. However, it is important to note that while there is a correlation between sequence similarity and PCA-derived Euclidean distance, the model’s predictions are based on specific sequence motifs, not just overall sequence similarity as shown in Figure 4 and 6.
@@
Ablation study. We performed an ablation study on the 217 immunogenicity dataset to analyze design choices in FT-SelfPAD. Table 7 shows that incorporating shift and swap noise, along with re-initializing four pre-trained model layers before fine-tuning, improves performance. Notably, the model is less effective when trained from scratch or fine-tuned without random initialization of its four layers (see the first and second rows in Table 7).
Table 6: Comparing maximum and average similarity between the sequences in 217 dataset and the sequences at the centre of human sequences in the test set of PAD. TP/FP: True/False positive, TN/FN: True/False negative.
Similarity Chain Type TP FP FN TN
Maximum Light 91.74 89.25 79.33 75.21
Heavy 82.39 83.80 80.28 70.42
Average Light 56.50 61.75 55.56 53.59
Heavy 54.49 59.26 55.46 47.30
@@ 5 Conclusion
In this work, we present a framework, SelfPAD, to learn representations from the Patented Antibody Database, and demonstrate its usefulness by providing a robust model for humanness prediction - a practical challenge in therapeutic antibody development. SelfPAD opens new avenues for advancing the understanding and optimization of antibody sequences, ultimately facilitating the development of more effective and safe therapeutic interventions. Despite its advantages, one of the main drawbacks of our approach is that the PAD is a noisy dataset and further curation is needed to improve the signal-to-noise ratio. Moreover, the PAD introduces new biases, e.g., it biases towards lab optimised rather than natural sequences. Thus, we can also improve our results by incorporating data from other sources such as the OAS since it includes a large number of natural sequences. Finally, although this work focuses on humanness prediction, we can fine-tune SelfPAD for other downstream tasks such as predicting developability properties of antibodies.
Impact Statement
Data Privacy and Intellectual Property Rights. Our study involves the utilisation of patent data, which raises important considerations regarding data privacy and intellectual property rights. All the data used in our research have been obtained in compliance with the relevant regulations. We also point out that although sequences in the patent data are released and available in the public domain, their use is subject to intellecutal property rights and, thus, they cannot be used directly as a treatment for a particular disease/indication. Finally, the legal landscape regarding copyright and scientific data may evolve, so it is important to stay informed about the developments in this area.
Potential Biases in Patent Databases. By developing predictive models for antibody humanness, we aim to enhance the efficiency and safety of therapeutic antibody development. However, we acknowledge the broader ethical implications of our work and the possibility of biases inherent in patent databases, including selection bias and geographic bias, which may influence the representativeness of the data. The patent dataset includes both natural and engineered antibody sequences. However, the quantity of natural sequences is limited, which might impact the overall performance of the model. One approach to improve the robustness and generalisability of our findings would be to enrich the dataset by including more natural sequences from the Observed Antibody Space (OAS dataset) and other sources.
Limitations and Uncertainties. There are several constraints inherent in using patent data for predictive modelling. These include the inherent biases in the patent filings, the lack of standardised annotations for antibody sequences, and potential inaccuracies in patent databases. Additionally, identifiers of function (potential targets) in the patent dataset are noisy and may not be completely reliable. This is often due to a tendency toward risk aversion in safeguarding intellectual property, potentially resulting in patent filings containing sequences with varying functions. This can potentially corrupt the representation learned during our pre-training stage. We try to make the model robust to noise by employing various techniques as outlined in our response to previous question(s). Finally, we encourage future research to address these limitations through better data collection methods, validation studies, and robustness assessments.
Acknowledgements
We thank the anonymous reviewers for their constructive feedback, and appreciate the support from everyone in the MLAB program at AstraZeneca. Special thanks to Massimo Sammito, Owen Vickery, Sridhar Neelamraju, and Isabelle Sermadiras for valuable discussions.